According to one survey in India, 75% of Instagram users love REELS. Suppose that 25
Instagram users (randomly selected) have been approached in the university located in vile
parle. They have been asked about their status of like/ dislike the Instagram- REELS.
a) What is the probability that Exactly 15 of them would agree with the claim (or said they
love Insta-REELS)?
b) What is the probability that Exactly 20 of them would agree with the claim (or said they
love Insta-REELS)?
(5 Marks)
3.b. ‘Bhartdarshan’ is an Internet-based travel agency wherein customers can see videos of the
cities they plan to visit. The number of hits daily is a normally distributed random variable
with a mean of 10,000 and a standard deviation of 2,400.
a. What is the probability of getting more than 12,000 hits?
b. What is the probability of getting fewer than 9,000 hits?

Answer by Boreal(15235)

Exactly 15 would be 25C15*0.75^15*0.25^10 using binomial formula.
=0.0417
-
Exactly 20 would be 25C20*0.75^20*0.25^5=0.1645
-
z=(x-mean)sd,
>2000/2400 or 0.83
probability is 0.2033 that z>0.83
-
probability is for z<-5/12 (-1000/2400)
that is 0.3385